The current price of Estelle Corporation stock is $ 25.00. In each of the next two​ years, this stock price will either go up by 23 % or go down by 23 %. The stock pays no dividends. The​ one-year risk-free interest rate is 5.3 % and will remain constant. Using the Binomial​ Model, calculate the price of a​ one-year put option on Estelle stock with a strike price of $ 25.00.
Answer #1
High price=Current price*up move=25*1.23=30.75
Low price=Current price*down move=25*0.77=19.25
Risk neutral probability for up move
q = (e^(-risk free rate*time)-D)/(U-D)
=(e^(-0.053*1)-0.77)/(1.23-0.77)=0.38778
Put option payoff at high price (payoff H)
=Max(Strike price-High price,0)
=Max(25-30.75,0)
=Max(-5.75,0)
=0
Put option payoff at low price (Payoff L)
=Max(Strike price-low price,0)
=Max(25-19.25,0)
=Max(5.75,0)
=5.75
Price of Put option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.053*1)*(0.387783*0+(1-0.387783)*5.75)
=3.34